Write the expression as the sine or cosine of an angle.
cos5xcos7x-sin5xsin7x

Answer by ikleyn(52814)

.
. . . = cos(5x + 7x) = cos(12x).     ANSWER



Use the addition formula for cosine


    cos(a + b) = cos(a)*cos(b) - sin(a)*sin(b).


It is one of fundamental/basic formulas of trigonometry.

----------------

On this formula, see the lessons
    - Addition and subtraction formulas
    - Addition and subtraction formulas - Examples
in this site.